# IFT-UAM/CSIC-25-21

### v2

@@ Abstract
Inverse fluid solutions for $R$–symmetry breaking – The hydrodynamics of inverse PTs was presented for the first time in Ref. [83, 84]. There exist five different possible expansion modes with negative bulk velocities [83, 96]: i) inverse detonations (weak and Chapman-Jouguet (CJ)), ii) inverse deflagrations (weak and CJ), and iii) inverse hybrids (see Fig.5). The inverse detonation (see left panel of Fig.5) is obtained by glueing a reaction front with $\xi_{w}=v_{-}<c_{s,-}$ and a rarefaction wave going from $v(\xi_{w}^{+})$ to 00 at $\xi=c_{s,+}$. In the plasma frame, the velocity $v(\xi_{w}^{+})=\mu(v_{-},v_{+})$ with $v_{-}<v_{+}$ is always negative, where $\mu(v_{-},v_{+})=(v_{-}-v_{+})/(1-v_{-}v_{+})$ is the Lorentz transformed velocity. Notice however that since $\xi$ is positive the bubble is actually expanding. Across the rarefaction wave, namely from $\xi=c_{s,+}$ to $\xi=\xi_{w}$, the pressure as well as the velocity decreases.
The inverse deflagrations (see right panel of Fig.5) are obtained by glueing a reaction front with $v_{+}=\xi_{w}>v_{J,\rm inv}$, where $v_{J,\rm inv}$ is the inverse Jouguet velocity [^6] This quantity, first introduced in [83], characterizes the transition between inverse detonations and inverse hybrid regimes and is defined as the velocity of the fastest (slowest) moving wall for an inverse detonation (inverse hybrid)., and $v_{-}>c_{s,-}$ followed by a compression wave ending with a shock front.
Finally, in the regime between inverse detonations and inverse deflagration, for $c_{s,+}>\xi_{w}>v_{J,\rm inv}$ the steady state is an inverse hybrid (see middle panel of Fig.5). This solution consists of a rarefaction wave glued to a detonation front, followed by a compression wave and finally a shock wave. We can also define the largest possible window in which inverse hybrid solutions can exist by selecting the slowest possible inverse Jouguet velocity. This constraint leads to the condition $c_{s,-}^{2}<\xi_{w}<c_{s,+}$ (see Appendix C for more details).
@@
This classification of hydrodynamic solutions was obtained within the (simplified) bag EoS. We have checked that this picture remains qualitatively the same also when considering the full form of the free energy (or effective potential) as evaluated explicitly for the SUSY model under consideration. In practice, we find only some quantitative differences related to the actual value of the speed of sound, which generally differs from $c_{s}^{2}=1/3$, and to the (mild) temperature dependence of $c_{s}^{2}$, which requires solving the coupled system of fluid equations for the pressure and the energy density as given in Eqs. (26) and (27). Explicit profiles obtained by solving numerically the fluid equations for the benchmark point with $\lambda=1.67$ and $m/\sqrt{F}=2$ are shown in detail in Fig. 5 for inverse detonations, inverse hybrids, and inverse deflagrations.
@@
Another possibility is that the bubble wall never reaches any of the steady states presented above and keeps accelerating until bubbles collide, namely it runs away. The terminal velocity $\xi_{w}$ of the bubble cannot be determined by hydrodynamic considerations only but requires a microscopic treatment. Employing the machinery presented in Ref. [83], we find in Appendix E that the bubble never runs away in the model that we study and always reaches one of the steady states. Our preliminary analysis however cannot determine which one of them.
Figure 5: Fluid velocity profiles for inverse phase transitions in the plasma frame, as evaluated by solving the hydrodynamic equations with the thermodynamics derived directly from the full free energy. Due to their inverse nature, the fluid velocity for these solutions is always negative or zero in this frame. From left to right, inverse detonation, hybrid, and deflagration transitions are displayed. Further details on the computation are provided in the main text and the Appendix C.
@@
Coupling to the SM thermal bath – In the early universe, the SUSY breaking model considered here is generally accompanied by additional spectator fields [^7] A spectator field does not change its physical properties, e.g. its mass, during the phase transition. that are in thermal equilibrium with the SUSY breaking sector. To assess the impact of these additional degrees of freedom, we redefine the energy and pressure as
@@
where we consider $\tilde{c}^{2}=1/3$ and $\tilde{a}$ controls the number of the relativistic spectator degrees of freedom (dofs), which is expected to be $\tilde{a}\sim 200\pi^{2}/30\sim 70$ considering a supersymmetric extension of the Standard Model (SM).
@@
$\frac{3}{4}|\alpha_{\vartheta}|<\frac{\omega_{-}(T_{+})-\omega_{+}(T_{+})}{\omega_{+}(T_{+})}=\frac{\Delta a(T_{+})}{a_{+}(T_{+})},$ (11)
where $a_{+}(T_{+})$ indicates the effective number of relativistic dofs in the symmetric phase at the temperature $T_{+}$, according to the parametrization of the enthalpy as $\omega/T^{4}\equiv a(T)$, and $\Delta a(T_{+})$ is the change in dofs in the broken phase at the same temperature. This relation indicates that an inverse FOPT can be very strong only when it involves a significant change in dofs between the two phases. This is a structural property of the vacua of the theory under consideration, and it should be contrasted with the case of standard FOPTs whose strength is mostly controlled by the amount of supercooling that can be achieved in the expanding universe. In particular, Eq. (11) indicates that an inverse FOPT is not necessarily stronger when it becomes more supercooled.
Conclusion and outlook – We presented a simple SUSY breaking model displaying a window of inverse FOPTs during the spontaneous breaking of the $R$–symmetry. This represents the first concrete example of a BSM model leading to an inverse FOPT in a cooling cosmology, as well as a proof of principle for the relevance of this dynamics in the early universe.
@@
As the actual hydrodynamic solution that will be realised depends on the terminal velocity of the bubble wall, our study suggests to apply and adapt the non–equilibrium techniques developed for direct FOPTs to inverse FOPTs as well.
@@
Appendix A Effective potential
@@
[…]
where the sum includes the (tree–level) massless fields in the $X$ superfield corresponding to the goldstino, the pseudomodulus, and the $R$–axion, which will give a constant, namely $x$–independent, contribution to the free energy at one loop.
@@
$V_{\rm eff}(x,T)=V_{\rm tree}(x)+\sum_{i}\left[V_{\rm CW}\big{(}M_{i}^{2}(x)\big{)}+V_{T}\big{(}M_{i}^{2}(x)\big{)}\right],$ (16)
where we have neglected the thermal masses as they play no significant role for the FOPT under study, given that the particles contributing to the effective potential are either very massive or very light in both phases.
Appendix B Nucleation temperature
@@
In this section we study the tunnelling rate for the $R$–symmetry breaking FOPT of interest. Starting with the effective potential derived in the previous section, we can proceed to the analysis of the phase transition. FOPTs occur when the minima of the effective potential corresponding to different phases are separated by a potential barrier, so that the transition proceeds via the nucleation of bubbles. The probability of bubble nucleation per unit time and unit volume is given by [100, 101, 102]
@@
where $S_{3}$ and $S_{4}$ are the $O(3)$ and $O(4)$ bounce actions, respectively, and $R_{0}$ is the bubble radius at nucleation. In the case of interest, the tunnelling is dominantly induced by thermal fluctuations, and we can thus neglect the $O(4)$ contribution. The probability of finding a specific point of the universe in the false vacuum at a given temperature is given by [103, 104]:
@@
Appendix C Solving the Hydrodynamic Equations for the Fluid Profiles
The conservation of the energy-momentum tensor for a relativistic fluid, given by
$\nabla_{\mu}T^{\mu\nu}=0,$ (23)
yields two independent hydrodynamic equations: one for energy conservation and another for momentum conservation. These equations can be rewritten in terms of the enthalpy density, $w=e+p$, as
$\partial_{\xi}\left[w\gamma^{2}v\right]$ $=0,$ (24)
$\partial_{\xi}\left[w\gamma^{2}v^{2}+p\right]$ $=0.$ (25)
We consider a spherically symmetric and self-similar configuration, where the fluid variables depend only on $\xi=r/t$, the similarity variable. In this variable, we can express Eqs. (24) and (25) in differential form
$(\xi-v)\frac{\partial_{\xi}e}{w}$ $=\frac{2v}{\xi}+[1-\gamma^{2}v(\xi-v)]\partial_{\xi}v,$ (26)
$(1-\xi v)\frac{\partial_{\xi}p}{w}$ $=\gamma^{2}(\xi-v)\partial_{\xi}v.$ (27)
To express Eq. (27) in terms of $v$ and $T$, we use the thermodynamic relation
$\frac{\partial_{\xi}p}{w}=\frac{\partial_{\xi}T}{w}\frac{dp}{dT}=\frac{\partial_{\xi}T}{T}.$ (28)
This leads to the temperature evolution equation
$\frac{\partial_{\xi}T}{T}=\gamma^{2}\mu(\xi,v)\partial_{\xi}v,\qquad\mu(\xi,v)=\frac{\xi-v}{1-\xi v}\,,$ (29)
where $\mu(\xi,v)$ is the Lorentz transformed velocity. Thus, the coupled system of hydrodynamic equations can be rewritten solely in terms of the velocity and temperature profiles, $v(\xi)$ and $T(\xi)$,
$(\xi-v)\frac{\partial_{\xi}T}{w}\frac{de}{dT}$ $=\frac{2v}{\xi}+[1-\gamma^{2}v(\xi-v)]\partial_{\xi}v,$ (30)
$\frac{\partial_{\xi}T}{T}$ $=\gamma^{2}\mu(\xi,v)\partial_{\xi}v.$ (31)
It is important to emphasize that the thermodynamic quantities, such as $p$ and $e$, must be evaluated in the appropriate phase depending on the region where the equation is being solved. Specifically, when computing the rarefaction wave of a detonation, the relevant phase is the newly formed one, corresponding to the true vacuum. Conversely, in the case of an inverse detonation, the quantities must be evaluated in the initial phase, corresponding to the false vacuum.
In the remainder of this section, we present the different types of expansion modes for inverse PTs within this general framework.
Inverse Deflagration
$\xi_{w}$ $=v_{+},\qquad v(\xi_{w})=\mu(v_{+},v_{-}),$ (32)
$T(\xi_{w}^{+})$ $=T_{+},\qquad T(\xi_{w}^{-})=T_{-},$ (33)
where the $+$ phase corresponds to the false vacuum, while the -- phase corresponds to the true vacuum.
Additionally, we impose the condition for the formation of a shock wave, which is given by
$\mu(\xi_{sh},v(\xi_{sh}))\xi_{sh}=c_{s,-}^{2}(T(\xi_{sh}))\,.$ (34)
Here we emphasize that, as we are considering the full temperature dependence of the free energy of our system rather than the approximate form based for instance on the bag EoS, the speed of sound is temperature dependent and must be evaluated in the -- phase at the temperature corresponding to the given point in the fluid profile. An example of such a solution is shown in the right panel of Fig. 5.
These initial conditions also apply to standard detonations, provided that the pair $(v_{+},v_{-})$ satisfies the necessary condition $v_{+}>v_{-}$.
Inverse Detonations
$\xi_{w}$ $=v_{-},\qquad v(\xi_{w})=\mu(v_{-},v_{+}),$ (35)
$T(\xi_{w}^{+})$ $=T_{+},\qquad T(\xi_{w}^{-})=T_{-}\,.$ (36)
It can be checked directly that the rarefaction wave terminates at
$\xi_{\rm end}=c_{s,+}(T(\xi_{\rm end}))\,.$ (37)
For a standard detonation, the substitution $c_{s,+}\to c_{s,-}$ must be applied, as the rarefaction wave develops behind the reaction front, i.e., in the new phase.
These initial conditions also apply to standard deflagrations, provided that the pair $(v_{+},v_{-})$ satisfies the appropriate conditions. In this case, the shock condition in Eq. (34) must be modified by replacing $c_{s,-}$ with $c_{s,+}$, as the shock forms ahead of the reaction front in the old phase.
Before discussing the last type of solution, it is important to highlight the presence of strong solutions in Fig. 7, where the red-shaded region indicates their domain. For (inverse) detonations, the strong regime is defined by the conditions
[…]
Similarly, for (inverse) deflagrations, we have
[…]
As previously discussed in [83], strong (inverse) detonations cannot be consistently realised, while strong (inverse) deflagrations, although they may initially form due to the dynamics of the phase transition, are inherently unstable. Over time, they will decay into (inverse) hybrid solutions.
Inverse Hybrid
$v(\xi_{w}^{+})$ $=\mu(\xi_{w}^{+},c_{s,+}(T_{+})),\qquad v(\xi_{w}^{-})=\mu(\xi_{w}^{-},v_{-}),$ (40)
$T(\xi_{w}^{+})$ $=T_{+},\qquad T(\xi_{w}^{-})=T_{-}\,,$ (41)
Additionally, the shock formation condition given by Eq. (34) must be imposed, and one can verify that the rarefaction wave of the inverse detonation terminates at Eq. (37). As explained in the main text, the maximal range of wall velocities for which an inverse hybrid solution exists is given by
$c_{s,-}^{2}<\xi_{w}<c_{s,+}\,,$ (42)
where the lower bound arises because the slowest possible inverse hybrid is determined by the slowest possible shock.
For the case of a direct hybrid transition, a strong deflagration must instead be connected to a CJ detonation, where the latter is characterized by $v_{-}=c_{s,-}(T_{-})$. The allowed range of wall velocities in this case is
$c_{s,-}<\xi_{w}<1\,,$ (43)
where the upper bound is simply the speed of light, as there is no fundamental constraint on the maximum speed of the shock front.
Overlap in the hybrid corner
Figure 7: Overlap of direct and inverse branches in the $(v_{-},v_{+})$ plane and corresponding fluid profiles. Left panel: The $(v_{-},v_{+})$ trajectories for different values of $T_{+}$. The inverse branch is shown in orange, while the direct branch is displayed in blue. The highlighted crossing point indicates a case where both a direct and an inverse solution exist for the same $(v_{-},v_{+})$ pair. Middle panel: Fluid profile corresponding to the direct hybrid solution. Right panel: Fluid profile for the inverse hybrid solution. The shaded regions indicate the interior of the bubble. The numerical values of $\alpha_{\vartheta}$ and relevant hydrodynamic parameters are reported in each panel.
In our numerical analysis, we observe that in the hybrid transition regime, the branches in the $(v_{-}-v_{+})$ plane exhibit an overlap between direct and inverse transitions. This is particularly evident when zooming in on the hybrid region, as shown in Fig. 7 (left panel). There, we explicitly construct two distinct solutions corresponding to the same pair of values $(v_{-},v_{+})$, demonstrating the existence of overlapping branches, in the middle and right panel of Fig. 7.
This overlap arises due to the stability conditions required for hybrid solutions. Specifically, for both direct and inverse hybrids to remain stable, the fluid velocity just behind (or in front of) the wall must match the local speed of sound in the respective phase at the corresponding temperature. That is, stability demands that
[…]
This condition provides additional flexibility in setting $\xi_{w}=v_{-}$ for direct hybrids and $\xi_{w}=v_{+}$ for inverse hybrids, thus allowing both solutions to coexist.
Another key reason for this overlap is related to the structure of the separatrices (black solid lines) in the $(v_{-},v_{+})$ plane. Ideally, these separatrices would be given by
$v_{-}=v_{+},\quad v_{-}v_{+}=c_{s,-}^{2},$ (45)
but since the speed of sound varies along the branches due to temperature dependence, the boundary between the direct and inverse solutions is no longer sharply defined.
Despite their overlap in the $(v_{-},v_{+})$ plane, the two solutions can still be distinguished physically. Each branch corresponds to a different set of temperatures $(T_{+},T_{-})$, leading to a different transition strength characterized by the generalised pseudotrace, $\alpha_{\vartheta}$, which will have in fact a different sign. Thus, even though the solutions may appear degenerate in velocity space, they remain distinct due to their thermodynamic properties. The matching conditions and the sign of $\alpha_{\vartheta}$ remain robust criteria for distinguishing direct and inverse transitions.
Appendix D Inverse FOPTs in $\mu\nu$ model
$e_{+}(T)=a_{+}T^{\nu}+\epsilon_{+}\,,\quad\qquad\ \ p_{+}(T)=c_{s,+}^{2}a_{+}T^{\nu}-\epsilon_{+}\,,$ (46)
$e_{-}(T)=a_{-}T^{\mu}+\epsilon_{-}\,,\quad\qquad\ \ p_{-}(T)=c_{s,-}^{2}a_{-}T^{\mu}-\epsilon_{-}\,,$ (47)
where the constants $\nu$, $\mu$ are related to the sound speed in the symmetric and broken phases through
$\nu=1+\frac{1}{c^{2}_{s,+}}\,,\quad\mu=1+\frac{1}{c^{2}_{s,-}}\,.$ (48)
The symmetric phase within our SUSY model corresponds to the $+$ phase, where $\langle x\rangle=0$, and we consider $\mu>\nu$ as this mimics the thermal history of the $R$–symmetry model. In this setup, the symmetric phase is energetically favoured both at $T=0$ and $T\gg T_{c}$, while the broken phase becomes dominant at intermediate temperatures.
In this model, the velocity relations from the matching conditions take the form:
$v_{+}v_{-}$ $=\frac{\mu-\mu\nu-r\nu(3\alpha_{\theta}-1)(\mu-1)}{\big{(}\mu-\mu\nu+r\nu(3\alpha_{\theta}+\mu-1)\big{)}(\mu-1)}\,,$ (49a)
$\frac{v_{+}}{v_{-}}$ $=\frac{(\mu-1)\big{(}\mu-\mu\nu+r\nu(3\alpha_{\theta}-1)\big{)}}{\mu-\mu\nu-r\nu(3\alpha_{\theta}+\mu-1)(\mu-1)}\,.$ (49b)
where we define the ratio $r$ as
$r\equiv\frac{a_{+}T^{\nu}_{+}}{a_{-}T^{\mu}_{-}}\,.$ (50)
Additionally, the strength parameter $\alpha_{\theta}$ defined from the pseudo trace $\theta$ as
$\theta=e-p/{c_{s,-}^{2}}\,\,,\quad\alpha_{\theta}\equiv\frac{4D\theta}{3w_{+}},$ (51)
within the $\mu\nu$ model evaluates to:
The direct (dashed lines) and inverse (solid lines) branches are presented in Fig. 8. As shown in the figure, as soon as $\alpha_{\theta}<0$, the inverse branches emerge. This confirms that in the $\mu\nu$-model, a negative $\alpha_{\theta}$ implies an inverse phase transition. Analogously, for the Bag EoS, a negative $\alpha_{+}$ corresponds to an inverse PT. This result aligns with the characterization proposed in [83], where it was shown that within the Bag EoS, $\Delta\epsilon<0$ serves as a direct indicator of an inverse phase transition.
Figure 8: Dashed lines represent direct phase transitions, while solid lines correspond to inverse phase transitions. The inverse branches emerge as soon as $\alpha_{\Theta}<0$, whereas this is not necessarily the case for $\alpha_{+}$. The two strength parameters of the phase transition, $\alpha_{+}$ and $\alpha_{\Theta}$, coincide in the bag model when $\mu=\nu=4$. The color gradient illustrates the variation of $\alpha_{\theta}$, highlighting the distinct regions of direct and inverse transitions.

### v3

@@ Abstract
Inverse fluid solutions for RR–symmetry breaking – The hydrodynamics of inverse PTs was presented for the first time in Ref. [26, 27] (see also [28]): there exist five different possible expansion modes with negative bulk velocities: i) inverse detonations (weak and Chapman-Jouguet (CJ)), ii) inverse deflagrations (weak and CJ), and iii) inverse hybrids.
@@
This classification of hydrodynamic solutions was obtained within the (simplified) bag EoS. We have checked that this picture remains qualitatively the same also when considering the full form of the free energy (or effective potential) as evaluated explicitly for the SUSY model under consideration. In practice, we find only some quantitative differences related to the actual value of the speed of sound, which generally differs from $c_{s}^{2}=1/3$, and to the (mild) temperature dependence of $c_{s}^{2}$, which requires solving the coupled system of fluid equations for the pressure and the energy density as discussed in Appendix  (B). An example of the explicit profiles obtained by solving numerically the fluid equations for the benchmark point with $\lambda=1.67$ and $m/\sqrt{F}=2$ are shown in detail in Fig. 3 for an inverse detonation, together with the free energy of the system at the nucleation temperature showing the direction of the phase transition and a sketch of the bubble with the corresponding fluid profile.
@@
Let us also mention that there is in principle the possibility that the bubble wall never reaches any of the steady states presented above, and keeps accelerating until bubbles collide, namely it runs away. Employing the line of reasoning presented in Ref. [26], we find that the bubble never runs away in the model under consideration, and always reaches one of the steady states (see App.,E and Refs. [5, 6, 7] for a derivation). Our hydrodynamic analysis however cannot determine which one of them, as mentioned above.
@@
Coupling to the SM thermal bath – In the early universe, the SUSY breaking sector considered here is generally accompanied by additional spectator fields  [^4] A spectator field does not change its physical properties, e.g. its mass, during the phase transition. that are in thermal equilibrium with the SUSY breaking sector and constitute a radiation bath. To assess the impact of these additional degrees of freedom, we redefine the energy and pressure as
@@
where $\tilde{c}^{2}=1/3$, and $\tilde{a}$ controls the number of the relativistic spectator degrees of freedom (dofs), which is expected to be $\tilde{a}\sim 70$ considering a supersymmetric extension of the Standard Model.
@@
$\frac{3}{4}|\alpha_{\vartheta}|<\frac{\omega_{-}(T_{+})-\omega_{+}(T_{+})}{\omega_{+}(T_{+})}=\frac{\Delta a_{\text{eff}}(T_{+})}{a_{{\text{eff}},+}(T_{+})}\,$, (9)
where $a_{{\text{eff}},+}(T_{+})$ indicates the effective number of relativistic dofs in the symmetric phase at the temperature $T_{+}$, according to the parametrization $3\omega(T)/4T^{4}\equiv a_{\text{eff}}(T)$, and $\Delta a_{\text{eff}}(T_{+})$ is the change in dofs in the broken phase at the same temperature. This relation shows that an inverse FOPT can be strong only when it involves a significant change in dofs between the two phases. This is a structural property of the vacua of the theory under consideration, and it should be contrasted with the case of standard FOPTs whose strength is mostly controlled by the amount of supercooling that can be achieved in the expanding universe. In particular, Eq. (9) indicates that an inverse FOPT is not necessarily stronger when it becomes more supercooled.
Conclusion and outlook – We presented a simple SUSY breaking model displaying a window of inverse FOPTs during the spontaneous breaking of the RR–symmetry. This represents the first explicit example of a BSM model leading to an inverse FOPT in a cooling cosmology, as well as a proof of principle for the relevance of this dynamics in the early universe.
@@
Appendix A Inverse FOPTs in the $\mu\nu$ model
$e_{\pm}(T)=a_{\pm}T^{\nu_{\pm}}+\epsilon_{\pm}\,,\quad\qquad\ \ p_{\pm}(T)=c_{s,+}^{2}a_{\pm}T^{\nu_{\pm}}-\epsilon_{\pm}\,,\quad\qquad\nu_{\pm}=1+1/{c^{2}_{s,\pm}}\,.$ (10)
In the following $\nu_{+}\equiv\nu$ and $\nu_{-}\equiv\mu$, we consider $\mu>\nu$ as this mimics the thermal history of the RR–symmetry model, presented in the main text. The velocity relations from the matching conditions take the form
$v_{+}v_{-}=\frac{\mu-\mu\nu-r\nu(3\alpha_{\theta}-1)(\mu-1)}{\big(\mu-\mu\nu+r\nu(3\alpha_{\theta}+\mu-1)\big)(\mu-1)}\,,\qquad\frac{v_{+}}{v_{-}}=\frac{(\mu-1)\big(\mu-\mu\nu+r\nu(3\alpha_{\theta}-1)\big)}{\mu-\mu\nu-r\nu(3\alpha_{\theta}+\mu-1)(\mu-1)}\,$, (11a)
where we define the ratio $r\equiv{a_{+}T^{\nu}_{+}}/{a_{-}T^{\mu}_{-}}$. Additionally, the strength parameter $\alpha_{\theta}$, defined from the pseudo-trace $\theta$ as $\alpha_{\theta}\equiv{4D\theta}/{3w_{+}}$ where $\theta=e-p/{c_{s,-}^{2}}$, within the $\mu\nu$ model evaluates to
It is shown in Fig. 4 that, as soon as $\alpha_{\theta}<0$, the inverse branches emerge. This confirms that in the $\mu\nu$-model, a negative $\alpha_{\theta}$ implies an inverse phase transition. Analogously, for the bag EoS, a negative $\alpha_{+}$ corresponds to an inverse PT. This result aligns with the characterization proposed in [26], where it was shown that within the bag EoS, $\Delta\epsilon<0$ serves as a direct indicator of an inverse phase transition.
Figure 4: Dashed (solid) lines represent direct (inverse) phase transitions. The inverse branches emerge as soon as $\alpha_{\theta}<0$, whereas this is not necessarily the case for $\alpha_{+}$. The two strength parameters of the phase transition, $\alpha_{+}$ and $\alpha_{\theta}$, coincide in the bag model when $\mu=\nu=4$.
Appendix B Solving the Hydrodynamic Equations for the Fluid Profiles
The conservation of the energy-momentum tensor for a relativistic fluid, given by $\nabla_{\mu}T^{\mu\nu}=0$, yields two independent hydrodynamic equations. These equations can be rewritten in terms of the enthalpy density, $w=e+p$. We consider a spherically symmetric and self-similar configuration, where the fluid variables depend only on $\xi\equiv r/t$, the similarity variable. Using this variable, it can be shown that the hydrodynamics equations, in terms of the fluid velocity $v(\xi)$ and the fluid temperature $T(\xi)$, take the following form
$(\xi-v)\frac{\partial_{\xi}T}{w}\frac{de}{dT}$ $=\frac{2v}{\xi}+[1-\gamma^{2}v(\xi-v)]\partial_{\xi}v\,,\qquad\frac{\partial_{\xi}T}{T}=\gamma^{2}\mu(\xi,v)\partial_{\xi}v\,$, (13)
where $\mu(\xi,v)=\frac{\xi-v}{1-\xi v}$. It is important to emphasize that the thermodynamic quantities, such as $p$ and $e$, must be evaluated in the appropriate phase depending on the region where the equation is being solved. In the remainder of this section, we present the different types of expansion modes for inverse PTs within this general framework.
Inverse Deflagration
$\xi_{w}=v_{+}\,,\qquad v(\xi_{w})=\mu(v_{+},v_{-})\,,\qquad T(\xi_{w}^{+})=T_{+}\,,\qquad T(\xi_{w}^{-})=T_{-}\,$, (14)
where the ++ phase corresponds to the false vacuum, while the - phase corresponds to the true vacuum. Additionally, we impose the condition for the formation of a shock wave, which is given by $\mu(\xi_{sh},v(\xi_{sh}))\xi_{sh}=c_{s,-}^{2}(T(\xi_{sh}))\,.$ These initial conditions also apply to standard detonations, provided that the pair $(v_{+},v_{-})$ satisfies the condition $v_{+}>v_{-}$.
Inverse Detonations
$\xi_{w}=v_{-}\,,\qquad v(\xi_{w})=\mu(v_{-},v_{+})\,,\qquad T(\xi_{w}^{+})=T_{+}\,,\qquad T(\xi_{w}^{-})=T_{-}\,.$ (15)
It can be checked directly that the rarefaction wave terminates at $\xi_{\rm end}=c_{s,+}(T(\xi_{\rm end}))$. For a standard detonation, the substitution $c_{s,+}\to c_{s,-}$ must be applied, as the rarefaction wave develops behind the reaction front, i.e., in the new phase.
These initial conditions also apply to standard deflagrations, provided that the pair $(v_{+},v_{-})$ satisfies the appropriate conditions. In this case, the shock condition must be modified by replacing $c_{s,-}$ with $c_{s,+}$, as the shock forms ahead of the reaction front in the old phase.
Before discussing the last type of solution, it is important to highlight the presence of strong solutions in Fig. 5, where the red-shaded region indicates their domain. For (inverse) detonations/deflagrations, the strong regime is defined by the conditions $(v_{+}\gtrless c_{s,+}(T_{+}))\ v_{-}\lessgtr c_{s,-}(T_{-})$. As previously discussed in [26], strong (inverse) detonations cannot be consistently realised, while strong (inverse) deflagrations, although they may initially form due to the dynamics of the phase transition, are inherently unstable. Over time, they will decay into (inverse) hybrid solutions.
Inverse Hybrid
$v(\xi_{w}^{+})=\mu(\xi_{w}^{+},c_{s,+}(T_{+}))\,,\qquad v(\xi_{w}^{-})=\mu(\xi_{w}^{-},v_{-})\,,\qquad T(\xi_{w}^{+})=T_{+}\,,\qquad T(\xi_{w}^{-})=T_{-}\,$, (16)
Additionally, the shock formation condition must be imposed, and one can verify that the rarefaction wave of the inverse detonation terminates again at $\xi_{\rm end}=c_{s,+}(T(\xi_{\rm end}))\,.$ The maximal range of wall velocities for which an inverse hybrid solution exists is given by $c_{s,-}^{2}<\xi_{w}<c_{s,+}$ where the lower bound arises because the slowest possible inverse hybrid is determined by the slowest possible shock.
For the case of a direct hybrid transition, a strong deflagration must instead be connected to a CJ detonation, where the latter is characterized by $v_{-}=c_{s,-}(T_{-})$. The allowed range of wall velocities in this case is $c_{s,-}<\xi_{w}<1$ where the upper bound is simply the speed of light, as there is no fundamental constraint on the maximum speed of the shock front.
Overlap in the hybrid corner
Figure 5: Overlap of direct and inverse branches in the $(v_{-},v_{+})$ plane and corresponding fluid profiles. Left panel: The $(v_{-},v_{+})$ trajectories for different values of $T_{+}$. The inverse branch is shown in orange, while the direct branch is displayed in blue. The highlighted crossing point indicates a case where both a direct and an inverse solution exist for the same $(v_{-},v_{+})$ pair. Middle panel: Fluid profile corresponding to the direct hybrid solution. Right panel: Fluid profile for the inverse hybrid solution. The shaded regions indicate the interior of the bubble.
In our numerical analysis, we observe that in the hybrid transition regime, the branches in the $(v_{-},v_{+})$ plane exhibit an overlap between direct and inverse transitions. This is particularly evident when zooming in on the hybrid region, as shown in Fig. 5 (left panel). There, we explicitly construct two distinct solutions corresponding to the same pair of values $(v_{-},v_{+})$, demonstrating the existence of overlapping branches, in the middle and right panel of Fig. 5.
This overlap arises due to the stability conditions required for hybrid solutions. Specifically, for both direct and inverse hybrids to remain stable, the fluid velocity just behind (or in front of) the wall must match the local speed of sound in the respective phase at the corresponding temperature. That is, stability demands that for (inverse) hybrid holds $(v_{+}=c_{s,+}(T_{+}))\ v_{-}=c_{s,-}(T_{-})$. This condition provides additional flexibility in setting $\xi_{w}=v_{-}$ for direct hybrids and $\xi_{w}=v_{+}$ for inverse hybrids, thus allowing both solutions to coexist.
Another key reason for this overlap is related to the structure of the separatrices (black solid lines) in the $(v_{-},v_{+})$ plane. Ideally, these separatrices would be given by $v_{-}=v_{+}$ and $v_{-}v_{+}=c_{s,-}^{2}$, however, since the speed of sound varies along the branches due to temperature dependence, the boundary between the direct and inverse solutions is no longer sharply defined. Despite their overlap in the $(v_{-},v_{+})$ plane, the two solutions can still be distinguished physically. Each branch corresponds to a different set of temperatures $(T_{+},T_{-})$, leading to a different transition strength characterized by the generalised pseudotrace, $\alpha_{\vartheta}$, which will have in fact a different sign. Thus, even though the solutions may appear degenerate in velocity space, they remain distinct due to their thermodynamic properties.
Appendix C Effective potential
@@
$J_{B/F}(y^{2})$ $=\int_{0}^{\infty}dx\ x^{2}\log\Big[1\mp\exp{(-\sqrt{x^{2}+y^{2}})}\Big]\,$, (20)
where the sum includes the (tree–level) massless fields in the $X$ superfield corresponding to the goldstino, the pseudomodulus, and the RR–axion, which will give a constant, namely xx–independent, contribution to the free energy at one loop.
@@
$V_{\rm eff}(x,T)=V_{\rm tree}(x)+\sum_{i}\left[V_{\rm CW}\big(M_{i}^{2}(x)\big)+V_{T}\big(M_{i}^{2}(x)\big)\right]\,.$ (21)
where we have neglected the thermal masses. In fact, perturbation theory for thermal field theory breaks down for massless particles. This breakdown can be avoided by the resummation of the so-called daisy diagrams, which amounts to add thermal masses to the tree-level ones in the CW and thermal potential. However, as our model does not contain dynamical massless particles on either phase, we expect this to be a good approximation. A detailed study of additional thermal effects is left for future work.
Appendix D Nucleation theory
@@
In this section we study the tunnelling rate for the RR–symmetry breaking FOPT of interest. Starting with the effective potential derived in the previous section, we can proceed to the analysis of the phase transition. FOPTs occur when the minima of the effective potential corresponding to different phases are separated by a potential barrier, so that the transition proceeds via the nucleation of bubbles. The probability of bubble nucleation per unit time and unit volume is given by [44, 90, 91]
@@
where $S_{3}$ and $S_{4}$ are the $O(3)$ and $O(4)$ bounce actions, respectively, and $R_{0}$ is the bubble radius at nucleation. In the case of interest, the tunnelling is dominantly induced by thermal fluctuations, and we can thus neglect the $O(4)$ contribution. The probability of finding a specific point of the universe in the false vacuum at a given temperature is given by [54, 65]
@@
For the FOPT under consideration, we find typical values of $\beta/H=\mathcal{O}(10^{4})$ in the relevant parameter space. Given the short duration of the phase transition, the corresponding GW signal will be suppressed. We leave a comprehensive study of the model parameter space in view of the detectability of current and future GW experiments for future work.